# KA-GNN: Kolmogorov-Arnold Graph Neural Networks for Molecular Property Prediction

### v1

@@ Abstract
Molecular property prediction is a crucial task in the process of Artificial Intelligence-Driven Drug Discovery (AIDD). The challenge of developing models that surpass traditional non-neural network methods continues to be a vibrant area of research. This paper presents a novel graph neural network model—the Kolmogorov-Arnold Network (KAN)-based Graph Neural Network (KA-GNN), which incorporates Fourier series, specifically designed for molecular property prediction. This model maintains the high interpretability characteristic of KAN methods while being extremely efficient in computational resource usage, making it an ideal choice for deployment in resource-constrained environments. Tested and validated on seven public datasets, KA-GNN has shown significant improvements in property predictions over the existing state-of-the-art (SOTA) benchmarks.
@@ 1 Introduction
Drug development is a complex and costly process, typically requiring decades of time and substantial investment [1, 2, 3]. In this challenging landscape, artificial intelligence (AI) has become particularly valuable, significantly impacting the prediction of molecular properties and showing immense promise in drug design [4, 5]. AI has greatly advanced virtual screening processes, potentially reducing the time and investment required [6, 7]. However, developing effective molecular representations and features remains a complex challenge. AI-based molecular models, which drive these advancements, generally fall into two categories: those based on molecular descriptors and end-to-end deep learning models [8].
@@
The first category relies on molecular descriptors or fingerprints as input features for machine learning algorithms. This process, known as featurization or feature engineering, involves not only capturing physical, chemical, and biological properties but also incorporating a wide array of fingerprints based on molecular structure information. Among these, structure-based fingerprints, particularly those derived from topological data analysis methods, have proven highly effective in molecular representation and featurization [9, 10, 11, 12]. The integration of these fingerprints with learning models has achieved significant success in various stages of drug design, including protein-ligand binding affinity prediction [13, 14, 15], protein mutation analysis [16, 17], and drug discovery [18], among other areas.
@@
The second category encompasses end-to-end deep learning models, where molecules are represented using Simplified Molecular Input Line Entry System (SMILES) strings, images, or molecular graphs, and various deep learning architectures such as Transformers, Convolutional Neural Networks (CNNs), and Graph Neural Networks (GNNs) are employed for molecular property prediction [19, 20, 21, 22]. Among these, molecular graphs, particularly those based on covalent bonds, are the most common representations of molecular topology at the atomic level. Numerous geometric deep learning models have been developed based on these molecular graphs, including Graph Convolutional Networks (GCNs) [23], graph autoencoders [24], and graph transformers [25]. These models are widely used in molecular data analysis and drug design. Additionally, molecular descriptors based on non-covalent interactions have recently demonstrated strong performance in predicting protein-ligand and protein-protein binding affinities [26, 12]. This suggests that new geometry-based molecular graph representations could potentially outperform traditional covalent-bond-based graphs. Integrating these geometry-based molecular graphs into geometric deep learning (GDL) models could enhance model performance and provide a better understanding of molecular geometry [27].
Figure 1: Overview of the KA-GNN model. A. Flowchart of the KA-GNN model. B. Example of a KAN at layer $l$ using Fourier series functions as pre-activation functions, illustrated for a case with 2 input dimensions and 3 output dimensions. C. Architecture for reconstructing node features used in the message-passing process of the GNN. D. Architecture of message passing in the KA-GNN at layer $l$.
Based on the Kolmogorov-Arnold representation theorem, Kolmogorov-Arnold Networks (KANs) have recently been proposed as a promising alternative to Multi-layer Perceptrons (MLPs) [28]. Due to their unique architecture and potential advantages, KANs stand out in several aspects. While MLPs, inspired by the universal approximation theorem, use fixed activation functions on nodes, KANs, inspired by the Kolmogorov-Arnold representation theorem, place learnable activation functions on edges, resulting in a fully-connected structure without linear weight matrices. This innovative design allows KANs to achieve higher accuracy and parameter efficiency compared to MLPs, particularly in solving partial differential equations. Recent research has explored various applications of KAN across different domains. Bozorgasl et al. [29] enhanced KAN by incorporating wavelet functions to improve interpretability and performance in capturing both high and low-frequency data components. Genet et al. [30] combined KANs with LSTMs, improving multi-step time series forecasting accuracy and efficiency. Cheon et al. [31] integrated KAN with pre-trained CNN models for remote sensing scene classification, demonstrating its effectiveness in this field.
In this paper, we present the development of KA-GNN, a novel model for molecular property prediction for the first time. The architecture of KA-GNN is shown in Figure 1. Our KA-GNN model leverages Fourier series and incorporates the Kolmogorov-Arnold Network (KAN) architecture to enhance the performance of a Graph Neural Network (GNN). In fact, our approach applies KAN to various components of the GNN, including node feature initialization, the message passing process, and graph classification tasks. A notable modification is the adjustment of the pre-activation function in KAN from B-splines to Fourier basis functions. Furthermore, our model considers both covalent bonds and non-covalent interactions defined by a cut-off distance, to construct the molecular graph representation in the GNN. These innovative adjustments ensure that KA-GNN remains parameter-efficient, while experimental results demonstrate that our model surpasses existing state-of-the-art pre-trained models across several benchmark tests. This highlights KA-GNN’s significant potential in drug development and molecular property prediction.
2 Related Work
In this section, we review the related research concerning molecular representation and then detail recent advances in graph neural networks for property prediction.
2.1 Kolmogorov-Arnold Network
Inspired by the Kolmogorov-Arnold representation theorem, Liu et al. [28] introduced the Kolmogorov-Arnold Network (KAN), a novel deep learning architecture that serves as a promising alternative to the traditional Multi-Layer Perceptron (MLP). Compared to MLP, KAN exhibits enhanced expressive power, parameter efficiency, high interpretability, and wide applicability.
Recent research has explored various applications of KAN in different domains. For instance, Bozorgasl et al. [29] incorporated wavelet functions into the Kolmogorov-Arnold network framework to improve interpretability and performance in capturing both high-frequency and low-frequency components of input data. Genet et al. [30] combined the strengths of KANs and LSTMs to enhance the accuracy and efficiency of multi-step time series forecasting. Additionally, Cheon et al. [31] integrated KAN with various pre-trained CNN models for remote sensing scene classification, demonstrating its efficacy in this domain.
These studies illustrate the versatility and potential of KAN in addressing diverse challenges across different fields, further validating its role as a robust and efficient deep learning architecture.
2.2 Molecular representation
Learning molecular representations is crucial for predicting the properties of compounds. At present, the modeling methods for molecular representations mainly fall into the following categories: sequence-based methods, graph-based methods and mixed data-based methods, [32, 33, 34]. Sequence-based methods[35, 36, 37] involve the use of sequences, such as SMILES strings, to represent molecules. These sequences are then processed using techniques inspired by natural language processing (NLP), taking advantage of the sequential nature of the data. However, the deep learning models are hindered by the nonunique nature of the SMILES string [36]. Graph-based methods consider molecules as graphs where atoms are nodes and bonds are edges. This category utilizes graph neural networks (GNNs) to capture the topological structure of molecules, making it highly effective for studying molecular connectivity and properties [38, 39, 40, 41]. AGBT[42] combined the SMILES string and a structure graph to construct molecular representation. DGL-LifeSci[43] integrated seven graph based models for molecular property prediction and pretrained models.
2.3 Graph Neural Network for Molecular Property Prediction
Graph neural network models have played a pivotal role in molecular data analysis. Traditional GNN models represent molecules as the defactor covalent-bond-based molecular graphs, and use major GNN architectures to learn molecular properties [44, 45, 46, 47]. With the importance of non-covalent bonds, cutoff-distance-based molecular graph representations have been widely employed in GNN models, such as DimeNet [48], Mol-GDL [27], etc. Further, higher-order interactions (beyond pair-wise forces) has been explicitly incorporated into GNN models, including ALIGNN [46], GEM [47], GemNet [49], etc, by the consideration of bond angles, dihedral angles, torsion angles, and other local geometric information. In particular, these higher-order terms can be directly related to MD force field information [50]. Finally, pre-training process has been adopted to further improve the accuracy of GNN models, such as N-Gram [51], PretrainGNN [52], GraphMVP[53], GEM [47], MolCLR [54], Uni-mol[55], SMPT[56], etc.
3 Method
3.1 Kolmogorov-Arnold representation theorem
@@
The Kolmogorov–Arnold Representation Theorem (or Superposition Theorem) is a milestone in the field of real analysis and approximation theory. It states that every multivariate continuous function can be represented as superposition of the addition of continuous functions of one variable. This theorem not only solves Hilbert’s thirteenth problem itself but also generalizes it to a broader form.
@@
3.2 Kolmogorov-Arnold Network (KAN)
@@
  • Although many constructive proofs of the Kolmogorov-Arnold representation theorem indicate that the inner functions $\phi_{q,p}$ in equation 1 are highly non-smooth [58, 59, 60], KAN opts for smooth functions as pre-activation functions $\phi_{j,i}^{(l)}$ to facilitate backpropagation. Liu et al. [28] selected functions based on B-splines. In this paper, we consider Fourier series functions.
Features Type Description Type Size
Atom CGCNN Atomic number, Radius, and electronegativity One-Hot 92
Bond Type Single, Double, Triple, or Aromatic. One-Hot 4
Bond Length Numerical and square length of the bond. Float 2
In Ring Indicates if the bond is part of a chemical ring. One-Hot 2
Cutoff Bond Atom charges Atoms charges in Molecular ($q_{i}$, $q_{j}$, $q_{i}$$\cdot$ $q_{j}$) Float 3
Table 1: Features Type and Description in KA-GNN model. Here we list both atom, covalent bond, and cuttoff bond features.
3.3 Graph Representation for Molecules
Molecules can naturally be viewed as graph structures, where the nodes represent atoms and the edges represent covalent bonds. Traditional molecular graphs typically include only the covalent bonds between atoms. However, in molecular dynamics, non-covalent interactions between atoms also significantly influence molecular properties. Therefore, when constructing the molecular graph, KA-GNN includes both covalent bonds and non-covalent interactions where atomic distances are less than 5 angstroms. These non-covalent bonds are treated as cut-off edges, providing a comprehensive reflection of the complex interactions within the molecule. The Graph Representation in the Figure 1 A illustrates the complex interactions within the molecule, highlighting both the covalent bonds (solid lines) and non-covalent cut-off bonds (dashed lines) with distances less than 5 angstroms are considered in our KA-GNN model.
The features of the graph are meticulously designed to encapsulate the various atomic properties and interactions detailed in molecular dynamics (MD) force fields. Specifically, our atomic features—comprising atomic number, radius, and electronegativity—are derived using Rdkit, following the approach in CGCNN [61]. Table 1 presents a comprehensive listing of the features for covalent bonds and non-covalent cutoff edges.
3.4 KA-GNN Method
Fourier-based KAN function
To optimize the network and avoid complex calculations, we utilize Fourier series as the pre-activation functions in Equation 2:
@@ 3.4 KA-GNN Method
Consequently, the activation value at the $j$-th neuron in layer $l+1$ can be obtained by:
@@
where $\mathbf{x}^{(l)}=(x_{1}^{(l)},x_{2}^{(l)},\dots,x_{n_{l}}^{(l)})$ denotes the input vector of activation values in layer $l$ of the KAN, and $\mathbf{x}^{(l+1)}=(x_{1}^{(l+1)},x_{2}^{(l+1)},\dots,x_{n_{l+1}}^{(l+1)})$ represents the output vector. Equation 4 can be concisely expressed as: $\mathbf{x}^{(l+1)}=\text{KAN}_{l}(\mathbf{x}^{(l)})$, where $\text{KAN}_{l}(\cdot)$ denotes the above KAN function in layer $l$. This network is integrated into a graph neural network, culminating in a novel architecture named KA-GNN, which is employed for molecular property prediction tasks.
Given that the inner functions in Equation 1 of the Kolmogorov-Arnold representation theorem can exhibit significant non-smoothness, we do not use this theorem as the foundational theory for showing the approximation capability of our model. Instead, we base our approach on the extension of Carleson’s theorem [62] regarding the convergence of Fourier series for multivariable functions, as proved by Fefferman in [63]:
@@
In summary, our model extends the Kolmogorov-Arnold Network (KAN) architecture by incorporating Fourier series as pre-activation functions to enhance approximation capability. By integrating this Fourier-based KAN within a Graph Neural Network (GNN) framework, we achieve a model with good performance and robustness.
@@
The model represents a molecule as a graph $G=(V,E)$, where $V$ denotes the set of nodes (atoms) and $E$ denotes the set of edges, defined as $E=E_{\text{chem}}\cup E_{\text{cutoff}}$. Here, $E_{\text{chem}}$ includes the covalent bonds, while $E_{\text{cutoff}}=\{uv\mid d(u,v)\leq 5\text{\AA}\}$ represents pairs of nodes within a cut-off distance of 5 angstroms. Each node $v\in V$ is associated with a feature vector $f_{v}$, which is a 92-dimensional vector composed of one-hot encoded representations of atomic properties, following the approach described in [61]. Similarly, each edge $uv\in E$ is associated with a feature vector $f_{uv}$, which is a 21-dimensional vector incorporating both chemical and geometrical information of the bond $uv$. Detailed descriptions of the feature vectors are provided in Table 1. The above feature vectors are utilized to reconstruct the initial node features $h_{v}^{(0)},v\in V$ for the graph neural network (GNN) in our model.
KA-GNN Model
the message passing are as following:
$h_{v}^{(0)}:=\text{KAN}_{\text{ini}}\left(f_{v}\oplus(\frac{1}{|N(v)|}\sum_{u\in N(v)}f_{uv})\right),$ (5)
where $N(v)$ is the set of neighbors of node $v$ ($v$ is not included), $|N(v)|$ denotes the number of neighbors of $v$, and $\oplus$ denotes the concatenation of vectors. The message passing process in layer $l$ can be expressed as follows:
$h_{v}^{(l+1)}:=h_{v}^{(l)}+\text{KAN}_{l}\left(h_{v}^{(l)}\oplus(\frac{1}{|N(v)|}\sum_{u\in N(v)}h_{u}^{(l)})\right).$ (6)
The initial KAN layer, $\text{KAN}_{\text{ini}}$, and the $l$-th KAN layer, $\text{KAN}_{l}$, correspond to the KAN function introduced in Equation 4, with their output dimensions set to 64. The input dimensions of these layers are adapted to fit the respective input vectors.
KA-GAT Model
by combining the characteristics of the attention score the message passing for the KA-GAT as as following:
The Mol-GAT utilized the graph representation features to generate the initial features $h_{v}^{(0)}$ and $h_{uv}^{(0)}$.

### v2

@@ Abstract
As key models in geometric deep learning, graph neural networks have demonstrated enormous power in molecular data analysis. Recently, a specially-designed learning scheme, known as Kolmogorov-Arnold Network (KAN), shows unique potential for the improvement of model accuracy, efficiency, and explainability. Here we propose the first non-trivial Kolmogorov-Arnold Network-based Graph Neural Networks (KA-GNNs), including KAN-based graph convolutional networks(KA-GCN) and KAN-based graph attention network (KA-GAT). The essential idea is to utilizes KAN’s unique power to optimize GNN architectures at three major levels, including node embedding, message passing, and readout. Further, with the strong approximation capability of Fourier series, we develop Fourier series-based KAN model and provide a rigorous mathematical prove of the robust approximation capability of this Fourier KAN architecture. To validate our KA-GNNs, we consider seven most-widely-used benchmark datasets for molecular property prediction and extensively compare with existing state-of-the-art models. It has been found that our KA-GNNs can outperform traditional GNN models. More importantly, our Fourier KAN module can not only increase the model accuracy but also reduce the computational time. This work not only highlights the great power of KA-GNNs in molecular property prediction but also provides a novel geometric deep learning framework for the general non-Euclidean data analysis.
Keywords Kolmogorov-Arnold Network, Fourier series, Graph Neural Network, Molecular Property Prediction
@@ 1 Introduction
Drug development is a complex and costly process, typically requiring decades of time and substantial investment [1]. In this challenging landscape, artificial intelligence (AI) has become particularly valuable, significantly impacting the prediction of molecular properties and showing immense promise in drug design [2, 3]. AI has greatly advanced virtual screening processes, potentially reducing the time and investment required [4, 5, 6]. AI-based molecular models, which drive these advancements, generally fall into two categories: those based on molecular descriptors and end-to-end deep learning models [7].
@@
The first category relies on molecular descriptors or fingerprints as input features for machine learning algorithms. This process, known as featurization or feature engineering, involves not only capturing physical, chemical, and biological properties but also incorporating a wide array of fingerprints based on molecular structure information. Among these, structure-based fingerprints, particularly those derived from topological data analysis methods, have proven highly effective in molecular representation and featurization [8, 9, 10, 11]. The integration of these fingerprints with learning models has achieved significant success in various stages of drug design, including protein-ligand binding affinity prediction [12, 13], protein mutation analysis [14, 15], and drug design [16, 17], among other areas.
@@
The second category includes end-to-end deep learning models that utilize various molecular representations such as Simplified Molecular Input Line Entry System (SMILES) strings, images, or molecular graphs, and deploy architectures like Transformers, Convolutional Neural Networks (CNNs), and Graph Neural Networks (GNNs) for prediction [18, 19, 20, 21, 22, 23]. Among these, molecular graphs based on covalent bonds are predominantly employed to describe molecular topology. Geometric deep learning models based on these molecular graphs, such as Graph Convolutional Networks (GCNs) [24], graph autoencoders [25], and graph transformers [26], have been extensively used in molecular data analysis and drug design. Additionally, recent research has demonstrated that molecular descriptors based on non-covalent interactions perform exceptionally well in predicting protein-ligand and protein-protein binding affinities [27]. These observations imply that new geometric molecular graph representations could surpass traditional covalent-bond-based graphs. By integrating these geometry-based molecular graphs into Geometric Deep Learning (GDL) models, it is possible to enhance model performance and deepen the understanding of molecular geometry [28].
Figure 1: Overview of the KA-GNN model architecture. The flowchart illustrates the modified components within the GNN: node embedding, message-passing, pooling and prediction modules.
Kolmogorov-Arnold Networks (KANs), which are based on the Kolmogorov-Arnold representation theorem, are increasingly recognized as a potent alternative to Multi-layer Perceptrons (MLPs). KANs distinguished by their unique architecture that employs different learnable activation functions, eliminate traditional linear weight matrices and enhance model accuracy and efficiency, particularly in solving partial differential equations, as described by Liu et al. [29]. Recent research highlights the versatility and adaptability of KANs across various domains [30, 31, 32, 33]. One notable application is the integration of KANs with existing neural network models to enhance performance. For instance, Genet et al. [33] significantly improved multi-step time series forecasting by integrating KANs with Long Short-Term Memory networks (LSTMs). Cheon et al. [34] effectively classified remote sensing scenes by merging KANs with pre-trained Convolutional Neural Network (CNN) models. Kiamari et al. [35] demonstrated their multifunctionality by incorporating KANs into Graph Convolutional Networks (GCNs) for semi-supervised graph learning tasks. Furthermore, adaptations in the base functions of KANs to better suit neural network applications have led to significant enhancements. Bozorgasl et al. [36] improved the interpretability and performance of KANs by incorporating wavelet functions that more effectively capture the frequency components of data. Tashin et al. [37] employed KANs with Fourier transform basis functions for feature transformation before GNN processing, validating their utility in Small Molecule-Protein Interaction Predictions. Li et al. [38] used adaptive Radial Basis Functions (RBFs) in KANs to enhance feature updating and replace MLPs in the prediction phase, demonstrating robust integration capabilities with various neural network frameworks. Kashefi et al. [39] applied Jacobi polynomials to design KAN layers for GNNs, effectively predicting fluid fields on irregular geometries. These advancements underscore the significant potential of KANs to refine neural network architectures. The ongoing exploration of how to further enhance KANs and integrate it into node feature embedding, message passing in different GNN frameworks, and the prediction phase remains a vital area of research.
In this paper, we introduce the first non-trivial Kolmogorov-Arnold Network-based Graph Neural Networks (KA-GNNs), including KAN-based convolutional networks (KA-GCN) and KAN-based graph attention network (KA-GAT). Figure 1 outlines the general KA-GNNs architecture. Different from all the previous trivial KAN-based GNN models, which only replace the MLP in the readout part with a standard KAN module, we utilizes KAN to optimize GNN architectures at three major levels, including node embedding, message passing, and readout. Further, a Fourier series-based KAN model has been developed and a rigorous theoretical prove of its robust approximation capability is given. Based on seven benchmark datasets, we have extensively validated our KA-GNNs and compared with state-of-the-art models. Our KA-GNNs have achieved both great accuracy and efficiency, providing a novel geometric deep learning framework for analyzing general non-Euclidean data.
2 Results
2.1 Kolmogorov-Arnold Network (KAN)
@@
Kolmogorov-Arnold Representation Theorem The Kolmogorov–Arnold Representation Theorem (or Superposition Theorem) is a milestone in the field of real analysis and approximation theory. It states that every multivariate continuous function can be represented as superposition of the addition of continuous functions of one variable. This theorem not only solves Hilbert’s thirteenth problem itself but also generalizes it to a broader form.
@@
2.1.1 Kolmogorov-Arnold Network (KAN)
@@
  • Although many constructive proofs of the Kolmogorov-Arnold representation theorem indicate that the inner functions $\phi_{q,p}$ in equation 1 are highly non-smooth [41, 42], KAN opts for smooth functions as pre-activation functions $\phi_{j,i}^{(l)}$ to facilitate backpropagation. Liu et al. [29] selected functions based on B-splines.
2.1.2 Fourier-series KAN model
To optimize the network and avoid complex calculations, we propose to utilize Fourier series as the pre-activation functions for KAN [37] as in Eq.(2):
@@ 2.1.2 Fourier-series KAN model
Consequently, the activation value at the $j$-th neuron in layer $l+1$ can be obtained by,
@@
where $\mathbf{x}^{(l)}=(x_{1}^{(l)},x_{2}^{(l)},\dots,x_{n_{l}}^{(l)})$ denotes the input vector of activation values in layer $l$ of the KAN, and $\mathbf{x}^{(l+1)}=(x_{1}^{(l+1)},x_{2}^{(l+1)},\dots,x_{n_{l+1}}^{(l+1)})$ represents the output vector. Eq (4) can be concisely expressed as: $\mathbf{x}^{(l+1)}=\text{KAN}_{l}(\mathbf{x}^{(l)})$, where $\text{KAN}_{l}(\cdot)$ denotes the above KAN function in layer $l$. This network is integrated into a graph neural network, culminating in a novel architecture named KA-GNNs, which is employed for molecular property prediction tasks.
Given that the inner functions in Eq (1) of the Kolmogorov-Arnold representation theorem can exhibit significant non-smoothness, we do not use this theorem as the foundational theory for showing the approximation capability of our model. Instead, we base our approach on the extension of Carleson’s theorem [43] regarding the convergence of Fourier series for multivariable functions, as proved by Fefferman in [44]:
@@
In summary, our model extends the Kolmogorov-Arnold Network (KAN) architecture by incorporating Fourier series as pre-activation functions to enhance approximation capability.
2.2 Kolmogorov-Arnold Network-based Graph Neural Network Models
Table 1: Features Type and Description in KA-GNN models. Here we list both atom, covalent bond, and non-covalent bond features.
Features Type Description Type Size
Atom CGCNN Atomic number, Radius, and electronegativity One-Hot 92
Bond Type Single, Double, Triple, or Aromatic. One-Hot 4
Bond Length Numerical and square length of the bond. Float 2
In Ring Indicates if the bond is part of a chemical ring. One-Hot 2
Non-Covalent Bond Atom charges Atoms charges in Molecular ($q_{i}$, $q_{j}$, $q_{i}$$\cdot$ $q_{j}$) Float 3
@@ 2.2 Kolmogorov-Arnold Network-based Graph Neural Network Models
In our KA-GNN models, a molecule is represented as a graph $G=(V,E)$, where $V$ denotes the set of nodes and $E$ denotes the set of edges. Each node represents an atom and an edge is formed among any two atoms if their distance is within a cutoff distance (in our model, we use cutoff distance as 5 Å). Figure 1 A illustrates the complex interactions within the molecule, highlighting both the covalent bonds (solid lines) and non-covalent cut-off bonds (dashed lines) with distances less than 5 angstroms are considered in our KA-GNN model. Our atomic features—comprising atomic number, radius, and electronegativity—are derived using Rdkit, following the approach in CGCNN [45] and PCNN [46]. Each node $v\in V$ is associated with a feature vector $f_{v}$, which is a 92-dimensional vector composed of one-hot encoded representations of atomic properties, following the approach described in [45]. Similarly, each edge $uv\in E$ is associated with a feature vector $f_{uv}$, which is a 21-dimensional vector incorporating both chemical and geometrical information of the bond $uv$. The edges in our model can be classified into two types, i.e., covalent bonds and non-covalent bonds, with different initial features. Detailed descriptions of the feature vectors are provided in Table 1.
2.2.1 KA-GCN Model
Our KA-GCN architecture has three major steps, including node embedding, message-passing, and readout. First, one KAN layer is employed in the initialization of node embeddings $h_{v}^{(0)},v\in V$ as follows,
$h_{v}^{(0)}:=\text{KAN}\left(f_{v}\oplus(\frac{1}{|N(v)|}\sum_{u\in N(v)}f_{uv})\right),$ (5)
where $N(v)$ represents the set of the neighboring nodes of node $v$ (excluding $v$ itself), $|N(v)|$ is the number of neighboring nodes, and $\oplus$ represents vector concatenation.
Second, one or several KAN layers are employed in the message passing module. At the layer $l$, the KAN-based message passing module can be expressed as,
$h_{v}^{(l+1)}:=h_{v}^{(l)}+\text{KAN}\left(h_{v}^{(l)}\oplus(\frac{1}{|N(v)|}\sum_{u\in N(v)}h_{u}^{(l)})\right).$ (6)
Computationally, the first KAN layer maps initial feature vector $f_{v}$ from dimension 113 to updated feature vector $h_{v}$ with dimension 64. After that, the feature dimension will stay the same in the later message-passing process.
Third, one or several KAN layers are employed in the readout module. After $L$-th the message passing phases, we apply average pooling across all node features to obtain a latent vector for the molecular graph $G$. An one-layer or two-layer KAN model is employed in readout module for the final prediction as follows,
$\hat{y}:=\text{KAN}\left(\frac{1}{|V|}\sum_{v\in V}h_{v}^{(L)}\right)\text{or},\hat{y}:=\text{KAN}\left(\text{KAN}\left(\frac{1}{|V|}\sum_{v\in V}h_{v}^{(L)}\right)\right)$
The cross-entropy is used as the loss function as follows,
$\mathcal{L}:=-\sum_{i}\left(y_{i}\log(\hat{y}_{i})+(1-y_{i})\log(1-\hat{y}_{i})\right),$
where $y_{i}$ denotes the actual label, and $\hat{y}_{i}$ denotes the predicted label.
2.2.2 KA-GAT Model
Similar to the KA-GCN architecture, the KA-GAT model initializes the node and edge embeddings, $h_{v}^{(0)}$ and $h_{uv}^{(0)}$, using a KAN layer as follows: